OpenStudy (akohl):

A 28-g bullet pierces a sand bag that is thick. The initial speed of the bullet is 55 m/s and it emerged from the sandbag at speed 18m/s What is the average magnitude of the friction force that slowed the bullet as it traveled through the bag? 1.3 N 130 N 38 N 13 N

OpenStudy (akohl):

Its s 30 cm thick

OpenStudy (abhisar):

use \(\large\sf v^2=u^2-2as\); Take v=18,u=55,s=30 `(convert in m frst)`

OpenStudy (abhisar):

find a

OpenStudy (akohl):

I keep getting 4501

OpenStudy (abhisar):

multiplied a with mass of bullet ?

OpenStudy (akohl):

When I multiply by mass I find friction is 126

OpenStudy (abhisar):

mass is in grams...convert in kg

OpenStudy (abhisar):

done ?

OpenStudy (akohl):

Thanks

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)